AO Commissioning on 15 October 2011

Observer(s):C. Arcidiacono, L. Busoni, A. Puglisi, L. Close, J. Males, JC. Guerra, K. Boutsia (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: T. Sargent (Tucson)

Telescope Support: J. Urban (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Clear sky when we open and winds <5m/s.

lightly cloudy.

Seeing about 0.7".

The system has been working with a high performances without major problems. The first half of the night has been dedicated to technical tests. -Change and test the Observe routines with the filter wheel changes and set the correct focus in the AO system. Filter vs focus. - Interaction Matrix on sky. Acquire a synthetic IM and sinusoidal IM. -Observing M15. Closing loop in crowd field. This has been a hard test for the AO system to close the loop with many stars and the autogain procedure didn't work well, at the end it was possible to close the loop. -Acquire and close the loop in faint star. Mag=16.5.

The neutral density filter on. -Observing ADI (HR9788) with ADI. Doing a quick reduction of the images shows clearly three planets. -Closing loop on the asteroid 45Eugenia. This requires a differemt approach to close the loop. The Observe has to use the Mode=track and the close loop manually. -Trapezoid. Observing using the observe routine without any problem. - H-K Tau. This observations has failed many times and it wasn't available to close the loop. Many problems to acquire the guide star and to build the catalogue -

Details

23:29 starting with some dome flats
  • we get error from pisces : ===> Error in pisces_send_request: unable to open connection
  • the server is down
  • server accessible again - start dome flat sequence in Ks.
  • durante gli ultimi dome flats the dome was open and there is light coming in and we were changing the filter

00:08 openining dome to start narrow band
  • start with Brg counts too low... ~1500 in 8sec
  • moving to FeII 7sec - still low - 2000counts
  • FeII 15sec - 4000counts
  • Ks 7sec - 10000counts - 9000counts
  • H 3sec - 11000counts
  • Ks 7sec - 8000counts
  • J 2sec - 12000counts
  • H 3sec - 10000counts
  • Ks 7sec - 6000counts
  • J 2sec - 9000counts
  • Ks 7sec - 4000counts
  • H 3sec - 4000counts
  • J 2sec - 4000counts

00:54 finished with flats

01:15 telescope is ready
  • AO group went up to check possible vignetting effect showed during the stages precision test.

1:30 telescope preset to the first field and collimation. Test of an upgrade of the Observe routine + AO pryramid focus procedure to set the right focus vs pisces filter.

1:53 Preset failed. Error message: star too bright for the AO.

02:14 Preset to M15. Testing closing the loop in a crowd field. Resend the same preset. 03:01 preset Active. Sky 3:21 Preset ACE in center M15

3:30 Preset failed. Active optics is not running. 3:31 Preset failed. Active optics problem. Maybe guide star too bright.

Change field and position of the telescope.

3:40 Preset AO786. (mag14).problem to find the guide star. Maybe too faint

3:56 Preset AO641. (mag:14.9) Failed to acquire the ref star. the background too bright and the ref. star shows mag:16.5

4:22 Preset AO750 (mag:15) ok. Images in open and close loop. Close loop at bin 4.

4:56 Preset HR8799. ADI observations. Really good results looking a quick reduction, the planets can be seeing and maybe we can see E, GREAT

07:13 Preset in TRACK mode trying to observe the asteroid "45 Eugenia" The close loop is send manually. Preset Ok. get_image failed because Pisces server died.
  • we finally found the asteroid
  • 4 offsets in J band - we see 'petit prince'
  • change to Ks - we clearly see the inner moon but we don't see 'petit prince' - not expected - probably will have to do longer integrations in Ks (now it is 1.5).
  • all offsets are set manually from the AO system. The observe routine thinks we are in open loop.
  • also some offsets in H band
  • probably the object over the asteroid in the J band was an artifact... (?).

08:44 Preset to HKtauB
  • preset, OBJECT=HKTauB, MODE=ace, GS=1, RS=0, PA=-90
  • we have no suitable guide stars around this target...
  • the preset keeps failing
  • we find a good star but the PA is not the correct one (pisces is upside down) - we need to cancel the preset
  • we find another GS in the correct PA but in the guider image the background is too high and the preset fails * we did not put the autocollimate command in the scripts but preset keeps failing

09:38 TO is checking pointing and collimation

09:43 We try to preset again at HKTauB
  • we give up on using the observe routine
  • we send a Track preset and will offset manually from the AO gui.
  • the telescope is way out of focus because we changed the arbitrator parameters.
  • we check pointing and collimation again on a bright star.
  • We try again the preset in Track mode and we collimate on the same star - we then remove the probe.
  • seeing 0.4''
  • the shell ripped - there are too few photons for some reason.
  • the position of the rotator is not passed from TCS to the AO... and we cannot close the loop.
  • we need to restart the SW.

10:54 The AO team goes to the dome to install the ND.

11:05 preset to TrapB
  • the guide stars are a bit bright but it is collimating
  • the J image is saturated (with the ND)... we try in FeII.
  • we practice the dither pattern using script TrapB_dither.obs - this worked

11:50 we change filter to Brg - so that specles can be identified - the AO group measure focus for Brg and then we repeat the dither pattern.

12:10 AO group does modulation tests using the TrapB as bright star target.

12:23 Preset to TrapC (GS=7) (PA=-35)
  • there are some clouds in the horizon
  • dither pattern in FeII - onels of the peaks of the binary are saturated
  • after the first pattern - the AO group decreases the Strehl to see if we manage to get an unsaturated image. - still saturated
  • change filter (Brg) and repeat dither. - last dither position is in open loop - it is too bright outside

13:15 - End of Night
  • we did not take ND flats because the sky is cloudy. We will try to take some tomorrow afternoon
  • we start dakrs for science and corquad ( observe,darkSET,/no_tcs )

Observe scripts comments.

  • Building the catalogue we found a scenario that the procedure gives error: Description. Editing the catalogue file to select only one object, the file simbad.txt has to be leave with a new line if the line is remove the procedure stop.
  • We are observing that the application pisces_image_display save the fits header with the information of the TCS and the AO information. This is a good tool when we are taking images from pisces. The only warning is that this procedure has to launch in another directory not used during the observations, eg:/20111015/HR8799/scratch/
  • When in the angle/PA command you press cancel, it keeps going with the preset
  • Increase filter wait time because the scirpt exits because of time out.

Comment.

It has been something about the way to archive the images that needs to be take care carefully. After setting the server on pisces to save the images, the images have been saving with the yesterday time stamp. This is a high risk that it will rewrite the yesterday images. The images processed by Observe in the running directory are ok, the raw data saved in /newdata are been rewritten. The yesterday images can be recovered from /warning directory and they have been saved in /OldRepository/backup-20111014/. A command "pisces_set_archive" is set. Now we have the right timestamp.

-- KonstantinaBoutsia - 14 Oct 2011
Topic revision: r12 - 31 Jan 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback